OpenStudy (anonymous):

Help Anyone? ammy has an annuity that pays him $9600 at the beginning of each year. Assume the economy will grow at a rate of 3.1% annually. What is the value of the annuity if he received it now instead of over a period of 10 years? $78,044.65 $80,651.16 $81,075.98 $83,999.29

OpenStudy (anonymous):

@hba help please?

OpenStudy (anonymous):

@amistre64 help please?

OpenStudy (anonymous):

@johnweldon1993 help please?

OpenStudy (anonymous):

@agent0smith help please?

OpenStudy (anonymous):

@ganeshie8 help please?

OpenStudy (amistre64):

not real sure how the growth rate of the economy factors in ....

OpenStudy (anonymous):

@ashwinjohn3

OpenStudy (anonymous):

@emilyowl

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

Sorry guys, his name is Sammy***

OpenStudy (anonymous):

Okay, so the easiest way I can explain it is complex. The problem is going to be 1) 9600 + (9600*3.1) = 29760 2) 29760 + (29760*3.1) = 92256 3) 92256 + (92256*3.1) = 378249.6 Once you get to the tenth step, divide it by ten.

OpenStudy (anonymous):

wait whats the 4rth step?

OpenStudy (agent0smith):

I'm not sure if you're meant to do it like compound interest, but i think maybe you are....